Determine the convergence interval of the equation below :

(x-5 neo nt3 20
Answer #1

(x-5) nt 3 Find anl n0o (2-5 Here an = n+3 (-5) an nco n3 -5 (n+3) -5 la-5 to convergu series For Raduu convergence -51 Set 2at endpans: Test Con (-) an \n+ 3 of form bn whch converges for 0ternatrig Serics Here, nc n+ 3 canverges So at an = n+3 wi t